# A unified linear intrinsic alignment model for elliptical and disc galaxies and the resulting ellipticity spectra

### v1

@@ Abstract
Alignments of disc galaxies were thought to result from tidal torquing, where tidal field of the cosmic large-scale structure exert torquing moments onto dark matter haloes, determining their angular momentum and ultimately the orientation of galactic discs. In this model, resulting intrinsic ellipticity correlations are typically present on small scales, but neither observations nor simulations have found empirical evidence; instead, simulations point at the possibility that alignments of disc galaxies follow a similar alignment model as elliptical galaxies, but with a weaker alignment amplitude. In our article we make the case for the theory of linear alignments resulting from tidal distortions of the galactic disc, investigate the physical properties of this model and derive the resulting angular ellipticity spectra, as they would appear as a contribution to weak gravitational lensing in surveys such as Euclid’s. We discuss in detail on the statistical and physical properties of tidally induced alignments in disc galaxies, as they are relevant for mitigation of alignment contamination in weak lensing data, comment on the consistency between the alignment amplitude in spiral and elliptical galaxies and finally, estimate their observability with a Euclid-like survey.
@@
thanks: ${}^{\star}$ basundharag@iisc.ac.inthanks: ${}^{\sharp}$ bjoern.malte.schaefer@uni-heidelberg.de
@@ 1 Introduction
The subject of galaxy alignments has been of prime importance and the correspondence of linear alignment models to elliptical galaxies and quadratic alignment models to spiral galaxies is well known (Schaefer(2009); Joachimi et al.(2015); Kiessling et al.(2015); Kirk et al.(2015); Troxel and Ishak(2014); Lamman et al.(2023)Lamman, Tsaprazi, Shi, Šarčević, Pyne, Legnani, and Ferreira). In general, any dark matter halo that contains galaxies undergoes a distortion due to gravitational tidal fields when embedded in the large-scale structure. The stellar component is oriented according to the perturbation of its dynamical equilibrium, and this orientation would depend on the alignment model being followed. If the tidal shear is small enough, the observed shape or the degree of ellipticity can be expected to have a linear relationship with the magnitude of the tidal field, described by a single alignment parameter (Hirata and Seljak(2004)). In the particular case of spiral galaxies, the galactic disc has an ellipticity that depends on its angle of inclination relative to the line of sight of the observer, which is determined by its angular momentum. Unlike in the case of elliptical galaxies where ellipticity is a measure of the shape and depends on the strength of the tidal shear, the ellipticity now is determined by the direction of the angular momentum. Although the relation between the angular momentum and the surrounding large-scale structure is not well-understood, one of the widely accepted mechanisms is that of tidal torquing, where even before a galaxy is formed, a torquing moment is exerted on the protohalo by the surrounding gravitational field (Crittenden et al.(2001)Crittenden, Natarajan, Pen, and Theuns). This imparts a default angular momentum to the host halo, and as the galactic disc forms later, it incorporates this angular momentum. In this kind of alignment, the tidal field in its lowest order enters in a quadratic form, hence the name quadratic alignment model. While linear tidal torquing theory is quite successful, the non-linear regime is not easy to describe analytically. The linear alignment model, on the other hand, is well-described even if the large-scale structure enters a nonlinear regime (nonlinear alignment model, see Hirata et al.(2007)Hirata, Mandelbaum, Ishak, Seljak, Nichol, Pimbblet, Ross, and Wake; Bridle and King(2007)). There have been studies to analyse the observability of intrinsic alignments of spiral galaxies. Crittenden et al.(2001)Crittenden, Natarajan, Pen, and Theuns; Schaefer and Merkel(2012); Chisari et al.(2015)Chisari, Codis, Laigle, Dubois, Pichon, Devriendt, Slyz, Miller, Gavazzi, and Benabed; Rodriguez-Gomez et al.(2017) derive ellipticity spectra for angular momentum-induced alignments. They should generate rather strong signals for Euclid-like (Amendola et al.(2018)) surveys (Schäfer and Merkel(2017)). Combining the alignment models for the two major galaxy types would lead to a rich phenomenology and the question of separating the two mechanisms (Tugendhat and Schäfer(2018); Reischke and Schäfer(2019)).
@@
Tidal shearing and tidal torquing have been recently tested in Zjupa et al.(2022)Zjupa, Schäfer, and Hahn; Delgado et al.(2023) using morphologically selected elliptical and spiral galaxies from the IllustrisTNG simulation (Nelson et al.(2018)) in the range of redshifts $10\leq z<$ 1. In this work it was found that elliptical galaxies show strong alignment signals further increasing with mass and redshift. Contrary to that, spiral galaxies exhibit a strong alignment only at high redshifts near $z=1$, and their alignment follows rather the tidal shearing instead of the tidal torquing model. Weak lensing surveys take a similar attitude, reporting a detection of a ellipticity cross-correlation between lensing and the shape of elliptical galaxies as a signature of tidal shearing. They place bounds on the cross-correlation of lensing and the shapes of spiral galaxies, implying that their alignment must be much weaker (Hildebrandt et al.(2017)Hildebrandt, Viola, Heymans, Joudaki, Kuijken, Blake, Erben, Joachimi, Klaes, Miller et al.; Johnston et al.(2019); Samuroff et al.(2023)), casting doubts on the validity of the quadratic alignment model.
@@
In this work, we propose a unified alignment model that can cater to both elliptical and spiral galaxies: If a such a model naturally accounts for the much weaker alignments of spirals, it would not be in conflict with observations and could account for the results from simulations. In particular, we investigate the possibility of a unified linear alignment model for disc galaxies, that follows an approach analogous to our computations for the ellipticity of elliptical galaxies in Ghosh et al.(2021)Ghosh, Durrer, and Schaefer. The underlying principle is that disc galaxies subjected to a gravitational tidal field undergo a distortion effect that can cause them to assume an elliptical shape, resulting in a change of the circular orbit of a star at the edge of a galaxy to an elliptical orbit. The star, which is bound by the NFW halo potential, experiences tidal forces, and the ellipticity of the galaxy scales with the tidal strength. An important ingredient for computing the ellipticity of galaxies is the alignment parameter which depends on the scale at which alignment due to tidal shear takes place. In case of elliptical galaxies, this scale can be assumed to be that of the Sérsic radius, following a de Vaucouleur’s profile with Sérsic index $n=4$, whereas for spiral galaxies, it should be typically the edge of the galactic disc. We perform a perturbative calculation to compute the effect of a quadrupolar tidal gravitational field on a circular orbit, which naturally generates an ellipticity proportional to the magnitude of the tidal field. We then simulate the orbit of a star in an NFW-halo, which is tidally perturbed again by a quadrupolar field, and this in turn is in agreement with our perturbative results. We eventually see that the strength of alignment for disc galaxies is smaller than that of elliptical galaxies. In following this idea, we want to provide a physical idea behind linear alignments for disc galaxies as observed in simulations (Hilbert et al.(2017)Hilbert, Xu, Schneider, Springel, Vogelsberger, and Hernquist; Jagvaral et al.(2022)Jagvaral, Singh, and Mandelbaum) and overpredicted amplitude of tidal torquing models which is in disagreement with observations (Heymans et al.(2013); Kilbinger et al.(2009)), as well as bridge to effective alignment models (Blazek et al.(2019)Blazek, MacCrann, Troxel, and Fang; Blazek et al.(2011)Blazek, McQuinn, and Seljak; Blazek et al.(2015)Blazek, Vlah, and Seljak). With a unified model for alignments of disc and elliptical galaxies, there would be no need to differentiate between the two types, leading to simplified effective models.
@@
Typically, we present results for a spatially flat standard $\Lambda$CDM-cosmology with conservative parameter choices: $\Omega_{m}=0.3$ (with a baryonic contribution of $\Omega_{b}=0.04$), $\Omega_{\Lambda}=1-\Omega_{m}$, $n_{s}=0.96$, $\sigma_{8}=0.8$ and a value of $h=0.7$. We assume an implicit summation over repeated indices. After introducing the linear alignment model for disc galaxies, investigating its physical properties and juxtaposing it with an analogous alignment model for elliptical galaxies in Sect. 2, we compute in Sect. 3 the resulting angular ellipticity spectra for a Euclid-like survey with a redshift distribution $n(z)\mathrm{d}z\propto z^{2}\exp(-z/z_{0})$ with a median redshift of 0.9. A summary and discussion in Sect. 4 concludes our investigation.
@@ 2.1 Scaling arguments
As the second moments of the brightness distributions are computed out to distances of $r_{\mathrm{vir}}$, the constant of proportionality, i.e. the alignment parameter $D_{\mathrm{IA}}$, has units of area and should be related to the squared size of the galaxy. One can compute dimensionless correction factors to this relation for brightness profiles of the Sérsic-family, resulting in small correction for exponential profiles, but rather large influences in the case of de Vaucouleurs-profiles (Ghosh et al.(2021)Ghosh, Durrer, and Schaefer; Giesel et al.(2022a)Giesel, Ghosh, and Schaefer). There should be a similar linear effect of tidal fields on the shape of a disc galaxy: The galactic disc extends typically to the NFW-scale radius $r_{s}$, which is a fraction $1/c(M_{\mathrm{vir}},z)$ of the virial radius $r_{\mathrm{vir}}$, giving rise to the phenomenon of flat rotation curves. The quantity $c(M_{\mathrm{vir}},z)$ is the concentration parameter which relates the virial mass $M_{\mathrm{vir}}$ of a halo to its redshift $z$. A quadrupolar tidal field distorts the edge of the galactic disc, compressing it in one direction and elongating it in the perpendicular direction, making it elliptical. The same scaling argument as in eqn. (1) suggests a relation of the form
@@ 2.2 Perturbative calculation
implying that the relative alignment strength in the same tidal field should be $\epsilon_{\mathrm{sp}}=t_{\mathrm{sp}}^{2}/t_{\mathrm{el}}^{2}\times\epsilon_{\mathrm{el}}$: The ratio between free-fall time scale and orbital time scale should determine the relative magnitude of the alignments.
@@ 2.3 Direct numerical simulation
$\frac{\mathrm{d}}{\mathrm{d}x}\Phi=\frac{GM_{\mathrm{vir}}}{Ar_{s}}\frac{x/(1+x)-\ln(1+x)}{x^{2}}$ (18)
@@ 2.4.1 Perturbed dynamics in disc and elliptical galaxies
Naively, the much smaller size $r_{s}$ of the galactic disc in comparison to the virial radius $r_{\mathrm{vir}}$ would be responsible for a much weaker alignment of disc galaxies relative to ellipticals. The alignment parameter for discs should be smaller by a factor of $1/c(M_{\mathrm{vir}},z)^{2}$ relative to the alignment of ellipticals, with $c$ being the NFW-concentration parameter. In fact, $r_{s}^{2}/r_{\mathrm{vir}}^{2}=1/c(M_{\mathrm{vir}},z)^{2}\simeq 0.02\ldots 0.06$, at redshifts around unity, where Euclid’s galaxy distribution is peaking, and for typical masses $10^{11\ldots 12}M_{\odot}/h$. The scaling of $c(M_{\mathrm{vir}},z)$ with redshift and halo mass follows the numerical model by Child et al.(2018)Child, Habib, Heitmann, Frontiere, Finkel, Pope, and Morozov, who condensed results from numerical simulations into phenomenological relations summarised by Fig. 3. On the other hand, the squared velocity ratio in the second term typically assumes values close to one, specifically $\sigma_{\mathrm{vir}}^{2}/\upsilon_{\mathrm{circ}}^{2}\simeq 0.9\ldots 0.95$ at a redshift of $z=1$ and for galaxy masses of $M=10^{11\ldots 12}M_{\odot}/h$ without a strong variability. Fig. 3 shows as well how the ratio between circular velocity $\upsilon_{\mathrm{circ}}$ at the scale radius $r_{s}$ and the velocity dispersion $\sigma_{\mathrm{vir}}^{2}$
@@
which can in principle lead to large numbers if $Q\simeq R$, as the integral diverges for this choice. Given the numerical results of Fig. 2, the averaging will yield larger values for $\langle\epsilon\rangle$ compared to the value for $\epsilon$ at $x=1$ by a few: $Q$ assumes values around unity, so that $R/Q\simeq 1/3$, and consequently, $\langle\epsilon\rangle\simeq 9/4=2.25$. Looking back at eqn. (21), this number corrects the expectation for $\epsilon_{\mathrm{sp}}$ in relation to $\epsilon_{\mathrm{el}}$ from $1/c(M_{\mathrm{vir}},z)^{2}\simeq 0.04$ to about 0.1, in agreement with the results by Hilbert et al.(2017)Hilbert, Xu, Schneider, Springel, Vogelsberger, and Hernquist found in a fluid-mechanical simulation of galaxy formation and evolution in terms of the relative amplitudes of the intrinsic alignment spectra separated by early- and late-type galaxies. More recently, simulations of Jagvaral et al.(2022)Jagvaral, Singh, and Mandelbaum find alignments of disc galaxies that are not as weak as those reported by Hilbert et al.(2017)Hilbert, Xu, Schneider, Springel, Vogelsberger, and Hernquist, which our model can accommodate by extending the radial averaging to larger distances, with a stronger response of the orbits to perturbing tidal fields. In summary, our analytical model does in fact explain ellipticities of disc galaxies that scale linearly with tidal field, and predicts the correct magnitude of the alignment. In the following, we work with a relative alignment parameter of $4$, intermediate between Hilbert et al.(2017)Hilbert, Xu, Schneider, Springel, Vogelsberger, and Hernquist and Jagvaral et al.(2022)Jagvaral, Singh, and Mandelbaum.

### v2

@@ Abstract
Alignments of spiral galaxies were thought to result from tidal torquing, where tidal field of the cosmic large-scale structure exert torquing moments onto dark matter haloes, determining their angular momentum and ultimately the orientation of galactic discs. In this model, resulting intrinsic ellipticity correlations are typically present on small scales, but neither observations nor simulations have found empirical evidence; instead, simulations point at the possibility that alignments of disc galaxies follow a similar alignment model as elliptical galaxies, but with a weaker alignment amplitude. In our article we make the case for the theory of linear alignments resulting from tidal distortions of the galactic disc, investigate the physical properties of this model and derive the resulting angular ellipticity spectra, as they would appear as a contribution to weak gravitational lensing in surveys such as Euclid’s. We discuss in detail on the statistical and physical properties of tidally induced alignments in disc galaxies, as they are relevant for mitigation of alignment contamination in weak lensing data, comment on the consistency between the alignment amplitude in spiral and elliptical galaxies and finally, estimate their observability with an Euclid-like survey.
@@
thanks: basundharag@iisc.ac.inthanks: bjoern.malte.schaefer@uni-heidelberg.de
@@ 1 Introduction
The subject of galaxy alignments is of prime importance for the interpretation of weak lensing data. In general, any dark matter halo that contains a galaxy undergoes tidal interactions due to gravitational tidal fields when embedded in the large-scale structure, and the stellar, luminous components becomes a proxy of this interaction, for instance by physical distortion of the halo away from its equilibrium shape, or by orientation, for instance relative to the halo’s angular momentum. These two alignment modes are discussed in the literature and have been conceived to describe alignments in elliptical (Hirata and Seljak, 2004; Hirata et al., 2007; Bridle and King, 2007) and spiral galaxies (Crittenden et al., 2001; Schaefer and Merkel, 2012; Chisari et al., 2015; Rodriguez-Gomez et al., 2017), respectively (Schaefer, 2009; Joachimi et al., 2015; Kiessling et al., 2015; Kirk et al., 2015; Troxel and Ishak, 2014; Lamman et al., 2023). Angular momentum-based alignments should generate rather strong ellipticity correlations and be easily measurable by Euclid-like (Amendola et al., 2018) surveys (Schäfer and Merkel, 2017). Combining the alignment models for the two major galaxy types would lead to a rich phenomenology and the question of separating the two mechanisms (Tugendhat and Schäfer, 2018; Reischke and Schäfer, 2019).
@@
Tidal shearing for ellipticals and tidal torquing for spirals have been recently tested in Zjupa et al. (2022); Delgado et al. (2023) using morphologically selected elliptical and spiral galaxies from the IllustrisTNG simulation (Nelson et al., 2018) in the range of redshifts $10\leq z<$ 1. In this work it was found that elliptical galaxies show strong alignment signals further increasing with mass and redshift. Contrary to that, spiral galaxies exhibit a strong alignment only at high redshifts near $z=1$, and their alignment follows rather the tidal shearing instead of the tidal torquing model. Weak lensing surveys take a similar attitude, reporting a detection of a ellipticity cross-correlation between lensing and the shape of elliptical galaxies as a signature of tidal shearing. They place bounds on the cross-correlation of lensing and the shapes of spiral galaxies, implying that their alignment must be much weaker (Hildebrandt et al., 2017; Johnston et al., 2019; Samuroff et al., 2023), casting doubts on the validity of the quadratic alignment model.
This is at par with the eagle simulation (Schaye et al., 2015; Crain et al., 2015; Velliscig et al., 2015) and the Illustris simulation (Vogelsberger et al., 2014a, b; Hilbert et al., 2017), whose results suggest that elliptical galaxies have a stronger alignment to that spiral galaxies of the same mass, which is also true for the host haloes of these galaxies. The MillenniumTNG simulation also finds that the alignment of spiral galaxies under a quadratic model is much weaker than that of elliptical galaxies under a linear model (Delgado et al., 2023). However, results from the MassiveBlack-II simulation (Tenneti et al., 2015) had shown that there is no significant difference between the direction of alignments in elliptical and spiral galaxies. Furthermore, Chisari et al. (2015) state, with respect to the horizon-AGN simulations, that the tangential orientation of disc galaxies around spheroidals are not safe to ignore in future weak lensing surveys. Some other simulations, like simba (Davé et al., 2019), emphasise that alignments of galaxies with respect to host filaments are largely dictated by their formation history and stellar mass function evolution. These wide range of results suggest that the discourse around galaxy alignments, especially the comparison between ellipticals and spirals, should be thoroughly pursued.
In this work, we propose a unified alignment model that can cater to both elliptical and spiral galaxies: If such a model naturally accounts for the much weaker alignments of spirals, it would not be in conflict with observations and could account for the results from simulations. It is to be noted that we do not claim our work to be the first instance where such an alignment is mentioned, and our work is mainly intended to outline a feasible microscopic mechanism for a unified alignment mechanism for both major galaxy types. In fact, previous works on perturbative models of galaxy alignments (Blazek et al., 2019; Vlah et al., 2020; Chen and Kokron, 2024) have already stated that both linear and higher-order terms exist simultaneously, and the alignment is predominantly linear or quadratic depending on the magnitude of the relative contribution from these terms. In itself, these effective field theories make the point that the dichotomy between tidal shearing and tidal torquing is too extreme. They assume that alignment models should contain the freedom to interpolate between these two bounding scenarios and state that a straightforward asssignment of alignment mechanisms to galaxy types is overly simplifying. In fact, Schaefer and Merkel (2012) discusses the case where the relative orientation between the tidal gravitational field and the moment of inertia of the galaxy determines the relative magnitudes of tidal shearing and tidal torquing within a halo.
Motivated by these results, we investigate the possibility of a unified linear alignment model for disc galaxies, that follows an approach analogous to our computations for the ellipticity of elliptical galaxies in Ghosh et al. (2021) and aim to establish a fundamental understanding of the respective alignment parameters. In the spirit of effective field theories, such a unified alignment model would imply that there is no need for separate alignment parameters for the two major galaxy types. Instead they would follow consistently from the internal dynamics of the galaxies and the characteristics of the stellar component. The underlying principle is that galactic discs subjected to a gravitational tidal field undergo a distortion effect that can cause them to assume an elliptical shape, resulting from the change of the otherwise circular orbit of a star in the galactic disc into an elliptical orbit. The ellipticity of the stellar orbits, which are bound by the Navarro-Frenk-White (NFW) halo potential (Navarro et al., 1996) should be linear in the external tidal shear.
@@
An important ingredient for computing the ellipticity of galaxies is the alignment parameter which depends on the scale at which alignment due to tidal shear takes place. In case of elliptical galaxies, this scale can be assumed to be that of the Sérsic radius, following a de Vaucouleur’s profile with Sérsic index $n=4$, whereas for spiral galaxies, it should be typically the scale radius of a galactic disc with typically exponential profile. We perform a perturbative calculation to compute the effect of a quadrupolar tidal gravitational field on a circular orbit, which naturally generates an ellipticity proportional to the magnitude of the tidal field. We then simulate the orbit of a star in an NFW-halo, which is tidally perturbed again by a quadrupolar field, and this in turn is in agreement with our perturbative results. We eventually see that the strength of alignment for disc galaxies is smaller than that of elliptical galaxies. In following this idea, we want to provide a physical idea behind linear alignments for disc galaxies as observed in simulations (Hilbert et al., 2017; Jagvaral et al., 2022) and overpredicted amplitude of tidal torquing models which is in disagreement with observations (Heymans et al., 2013; Kilbinger et al., 2009), as well as bridge to effective alignment models (Blazek et al., 2019, 2011, 2015). With a unified model for alignments of disc and elliptical galaxies, there would be no need to differentiate between the two types, leading to simplified effective models.
@@
Typically, we present results for a spatially flat standard $\Lambda$CDM-cosmology with conservative parameter choices: $\Omega_{m}=0.3$ (with a baryonic contribution of $\Omega_{b}=0.04$), $\Omega_{\Lambda}=1-\Omega_{m}$, $n_{s}=0.96$, $\sigma_{8}=0.8$ and a value of $h=0.7$. We assume an implicit summation over repeated indices. After introducing the linear alignment model for disc galaxies, investigating its physical properties and juxtaposing it with an analogous alignment model for elliptical galaxies in Sect. 2, we compute in Sect. 3 the resulting angular ellipticity spectra for an Euclid-like survey with a redshift distribution $n(z)\mathrm{d}z\propto z^{2}\exp(-z/z_{0})$ with a median redshift of 0.9. A summary and discussion in Sect. 4 concludes our investigation.
@@ 2.1 Scaling arguments
As the second moments of the brightness distributions are computed out to distances of $r_{\mathrm{vir}}$, the constant of proportionality, i.e. the alignment parameter $D_{\mathrm{IA}}$, has units of area and should be related to the squared size of the galaxy. One can compute dimensionless correction factors to this relation for brightness profiles of the Sérsic-family, resulting in small correction for exponential profiles, but rather large influences in the case of de Vaucouleurs-profiles (Ghosh et al., 2021; Giesel et al., 2022a). There should be a similar linear effect of tidal fields on the shape of a disc galaxy: The galactic disc extends typically to the NFW-scale radius $r_{s}$, which is a fraction $1/c(M_{\mathrm{vir}},z)$ of the virial radius $r_{\mathrm{vir}}$, giving rise to the phenomenon of flat rotation curves. The quantity $c(M_{\mathrm{vir}},z)$ is the concentration parameter which relates the virial mass $M_{\mathrm{vir}}$ of a halo to its redshift $z$. A quadrupolar tidal field distorts the edge of the galactic disc, compressing it in one direction and elongating it in the perpendicular direction, making it elliptical. The same scaling argument as in eqn. (1) suggests a relation of the form
@@ 2.2 Perturbative calculation
implying that the relative alignment strength in the same tidal field should be $\epsilon_{\mathrm{sp}}=t_{\mathrm{sp}}^{2}/t_{\mathrm{el}}^{2}\times\epsilon_{\mathrm{el}}$: The ratio between free-fall time scale and orbital time scale should determine the relative magnitude of the alignments. An interesting earlier work in this context is Camelio and Lombardi (2015), where the authors ascertain that an instantaneous external tidal field is not sufficient to explain the magnitude of the observed intrinsic alignment signals, and hence galaxy merger histories are also possible contributors to the signal. A discussion about galaxy evolution processes and their effect on our model is beyond the scope of our work, but will be an intriguing area to explore.
@@ 2.3 Direct numerical simulation
$\frac{\mathrm{d}\Phi}{\mathrm{d}x}=\frac{GM_{\mathrm{vir}}}{Ar_{s}}\frac{x/(1+x)-\ln(1+x)}{x^{2}}$ (18)
@@ 2.4.1 Perturbed dynamics in disc and elliptical galaxies
Naively, the much smaller size $r_{s}$ of the galactic disc in comparison to the virial radius $r_{\mathrm{vir}}$ would be responsible for a much weaker alignment of disc galaxies relative to ellipticals. The alignment parameter for discs should be smaller by a factor of $1/c(M_{\mathrm{vir}},z)^{2}$ relative to the alignment of ellipticals, with $c(M_{\mathrm{vir}},z)$ being the NFW-concentration parameter. In fact, $r_{s}^{2}/r_{\mathrm{vir}}^{2}=1/c(M_{\mathrm{vir}},z)^{2}\simeq 0.02\ldots 0.06$, at redshifts around unity, where Euclid’s galaxy distribution is peaking, and for typical masses $10^{11\ldots 12}M_{\odot}/h$. The scaling of $c(M_{\mathrm{vir}},z)$ with redshift and halo mass follows the numerical model by Child et al. (2018), who condensed results from numerical simulations into phenomenological relations summarised by Fig. 2. On the other hand, the squared velocity ratio in the second term typically assumes values close to one, specifically $\sigma_{\mathrm{vir}}^{2}/\upsilon_{\mathrm{circ}}^{2}\simeq 0.9\ldots 0.95$ at a redshift of $z=1$ and for galaxy masses of $M=10^{11\ldots 12}M_{\odot}/h$ without a strong variability. Fig. 2 shows as well how the ratio between circular velocity $\upsilon_{\mathrm{circ}}$ at the scale radius $r_{s}$ and the velocity dispersion $\sigma_{\mathrm{vir}}^{2}$
@@
Refer to captionRefer to caption
Figure 2: Scaling of the concentration parameter $c(M_{\mathrm{vir}},z)$ with virial halo mass $M_{\mathrm{vir}}$ and redshift $z$ for two common halo profile families, NFW (Navarro et al., 1996) and Einasto (Einasto, 1965) (top panel) and dependence of the squared velocity ratio $\sigma^{2}/\upsilon_{\mathrm{circ}}^{2}$ (bottom panel), both derived with the fitting formula from Child et al. (2018).
@@
which can in principle lead to large numbers if $Q\simeq R$, as the integral diverges for this choice. Given the numerical results of Fig. 3, the averaging will yield larger values for $\langle\epsilon\rangle$ compared to the value for $\epsilon$ at $x=1$ by a few: $Q$ assumes values around unity, so that $R/Q\simeq 1/3$, and consequently, $\langle\epsilon\rangle\simeq 9/4=2.25$. Looking back at eqn. (21), this number corrects the expectation for $\epsilon_{\mathrm{sp}}$ in relation to $\epsilon_{\mathrm{el}}$ from $1/c(M_{\mathrm{vir}},z)^{2}\simeq 0.04$ to about 0.1, in agreement with the results by Hilbert et al. (2017) found in a fluid-mechanical simulation of galaxy formation and evolution in terms of the relative amplitudes of the intrinsic alignment spectra separated by early- and late-type galaxies. More recently, simulations of Jagvaral et al. (2022) find alignments of disc galaxies that are not as weak as those reported by Hilbert et al. (2017), which our model can accommodate by extending the radial averaging to larger distances, with a stronger response of the orbits to perturbing tidal fields. In summary, our analytical model does in fact explain ellipticities of disc galaxies that scale linearly with tidal field, and predicts the correct magnitude of the alignment. In the following, we work with a relative alignment parameter of $4$, intermediate between Hilbert et al. (2017) and Jagvaral et al. (2022).